OpenStudy (brittanydosey):

You have a part-time job paying $9.95/hr and your deductions are FICA (7.65%), federal tax withholding (9.3%), and state tax withholding (5.8%). You save $30.00/wk for a savings account. If you work 15 hours/wk, how much discretionary income is left each week? Assume that all of the income from your part-time job is discretionary.

OpenStudy (brittanydosey):

Multiply your hourly wage by the number of hours that you work in a week. Then, using multiplication, remove the deductions. Lastly, subtract out the amount that you put into the savings account. ROUND TO THE NEAREST PENNY.

OpenStudy (brittanydosey):

@kelliegirl33

OpenStudy (anonymous):

9.95 * 15 = 149.25 deductions : FICA = .0765(149.25) = 11.417625 fed tax with = .093 (149.25) = 13.88025 state tax with = .058 (149.25) = 8.6565 total = 33.954375...rounds to 33.95 149.25 - 33.95 = 115.30 115.30 - 30 = $ 85.30

OpenStudy (brittanydosey):

thank you i have 2 more